OpenStudy (anonymous):

f(x)=(3x-3)e^-6x find the critical points

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=(3x-3)e^{-6x} }\)

OpenStudy (solomonzelman):

Did you find the derivative \(\large\color{black}{ \displaystyle f'(x) }\) ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{ d }{dx}\left[ ~A(x)\times S(x)~\right]=A'(x)S(x)+A(x)S'(x) }\)

OpenStudy (solomonzelman):

that is the product rule

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle \frac{ d }{dx}\left[ ~(3x-3)\times e^{-6x}~\right]=(3x-3)'\times e^{-6x}+(3x-3)\times (e^{-6x})' }\)

OpenStudy (anonymous):

ok I got that

OpenStudy (solomonzelman):

Ok, so the frst derivative f \('\)(x) is?

OpenStudy (anonymous):

that is the derivative

OpenStudy (solomonzelman):

(I want you to give me the equation of the derivative)

OpenStudy (anonymous):

\[3e ^{-6x}+(3x-3)e ^{-6x}\]